OpenStudy (asylum15):

Solve for 0 less than or equal to A less than or equal to 2pie 2sinA - 3cosA = 0

OpenStudy (anonymous):

2SinA=3CosA SinA/CosA=3/2 TanA=3/2 56.31=A

OpenStudy (anonymous):